# Improving Steering and Verification in AI-Assisted Data Analysis with Interactive Task Decomposition

### v1

@@ 6.3.3. Steering for Verification
[…]
@@ 6.3.4. Focusing on Smaller Steps Facilitated Verification
Additionally, participants experienced less information overload: “you get the blocks one-by-one so you are not overwhelmed by too much information” (P5), and compared to the Phasewise system which felt more like “debugging somebody else’s code than writing my own code” (P10). However, several participants (P4, P5, P13, and P18) were critical about how the Stepwise system discloses the next step. P18 stated that “I do not want it to do everything at once, but I want to know what it’s gonna do next” and P5 expressed reduced confidence for “not knowing the steps in advance.”
@@ 6.3.5. Aggregated Information Helped with Verification
[…]
@@ 6.3.7. Deferring Steering after Seeing Initial Results
Participants frequently preferred to first see results before interacting with and steering the AI tool. P6 highlighted that they “just want to see the result first and then trim it”. P18 wanted to “look at the result first and if the result was nonsense then go back”. This was particularly the case in the Phasewise and Stepwise systems, as P4 mentioned that the system made them “go through all of these steps and spend so much time before [they] could actually see the result.” P8 wished to see the generated code because they did not trust the system to generate correct code even if the execution plan was correct: “part of me wants to just see what it does, even if the [execution] plan looks reasonable, I know there’s gonna probably be errors.” Another reason why participants wanted to defer steering, particularly in the Phasewise system, was that the input/output assumptions or the execution plan did not have enough details about how the AI is going to “handle” or “calculate” something” (P10, P14). Similarly, P12 was not sure about something in the plan, so they said “I’m going to generate first and see what happens.” Lastly, P4 indicated that having a small working memory was why they preferred the baseline system, where every interaction with the AI resulted in a complete output to verify.
@@ 6.4. Summary of Results
Our study revealed that while there was no difference in task success, completion time, or number of required verification hints, participants felt significantly more in control of the data analysis process when using the Phasewise and Stepwise systems compared to the Conversational baseline. This was attributed to the ability to directly edit AI-generated assumptions and actions. However, the lack of immediate feedback after updating an assumption and the loss of downstream edits when updating upstream components were identified as limitations.
@@ 7. Discussion and Implications for AI-Assisted Data Analysis Tools
  • DQ2 Steering Support: How does the user determine which direction they should steer the AI? How should the tool facilitate users in making informed decisions at each steering point?
@@ 7.1. DQ1: Steering Points
Our results comparing user experiences with these systems reveals significant trade-offs. More frequent steering points increase users’ confidence in the results and their sense of control over the AI’s process. However, it demands more cognitive effort, delays the final outcome, and may lead to premature decisions as users commit to directions without understanding their future impact (in the terms of the Cognitive Dimensions of Notations framework (Green, 1989), this is a clear case of the system imposing “premature commitment”). Indeed, our findings indicate varied user preferences between the Conversational and Stepwise systems due to these factors.
@@ 7.2. DQ2: Steering Support
Participants noted their desire for the AI to act as an agent, aiding in the assumption-building process at steering points. Future tools could automatically retrieve assumptions by pinpointing specific, relevant evidence to support informed decision-making without overwhelming users. Additionally, these tools could help access to domain-specific knowledge pertinent to the data analysis task. Lastly, tools should be transparent regarding the AI’s limitations in sourcing all necessary information for optimal decision-making at each decision point.
@@ 7.3. DQ3: Steering Modality
After the user has decided which direction to steer the AI, their next action is to specify their intent to the AI. The decision question here is determining the right interface and modality for the user to steer the AI. In response, we introduced two distinct interfaces: a free-form text editor used in the Conversational system for maximal flexibility, and structured editors in the Stepwise and Phasewise systems. The structured editors contain the assumptions and actions generated by the LLM’s chain-of-thought, allowing users to edit them for steering.
@@ 8.2. System Limitations
Another limitation of the system is that edits to downstream code or assumptions are not propagated to upstream assumptions, and if a user makes an earlier edit it will overwrite any subsequent changes. The intention behind this prima facie design decision was to present a simple model of ”cause and effect” that represented how completions were generated, namely, the context of any point in the system is only that which appears before. Some participants identified this limitation and it influenced their steering preferences, choosing not to edit an assumption because it would clear changes that had been made to code.
@@ 9. Conclusion
In this work, we explore the design space of AI-assisted data analysis tools by presenting two novel interfaces that aim to improve steering and verification. Starting from the observation that task decomposition is an emerging characteristic of recent LLM-based systems, we developed two systems that explore different modes of interactive task decomposition, each based on unique trade-offs. The first, Stepwise, decomposes the problem step by step; the second, Phasewise, decomposes the problem by analysis logical phases. Our evaluation demonstrates that users experienced a greater sense of control and confidence with our systems in comparison to a chat-based baseline. Still, task decomposition is not without preference or cost. Some users prefer to work through the task incrementally, whilst others prefer to see the plan upfront. Additionally, highly-structured decomposition can introduce cognitive burden. Consequently, we imagine that future AI interfaces will need to support adaptive decomposition that reacts to the user and task.

### v2

@@ 6.3.3. Steering for Verification
Therefore, a recurring theme that emerged was participants trying to update the code, through steering, to include more informative and useful intermediary variables. For example, P16 added a new step to the execution plan to emit new outputs and other relevant columns in addition to just showing the final result. P4 added an explicit step to the execution plan display couple of groups so I can manually verify. Interestingly, there were also several cases that participants just did not understand the method used in the generated code, and therefore, asked the system to “come up with a more understandable solution” (P6).
@@ 6.3.4. Focusing on Smaller Steps Facilitated Verification
Additionally, participants experienced less information overload: “you get the blocks one-by-one so you are not overwhelmed by too much information” (P5), and compared to the Phasewise system which felt more like “debugging somebody else’s code than writing my own code” (P10). However, several participants (P4, P5, P13, and P18) were critical of the Stepwise system for not providing any information about the upcoming next step. P18 stated that “I do not want it to do everything at once, but I want to know what it’s gonna do next” and P5 expressed reduced confidence for “not knowing the steps in advance.”
@@ 6.3.5. Aggregated Information Helped with Verification
Many participants (n=7) appreciated how the Phasewise system aggregated descriptive statistics and assumptions for each column, finding that these elements scaffolded the AI’s reasoning about the task. These helped P15 “understand what are the different possibilities,” enabled P6 to “see exactly how each column will be treated”, and “forced” P3 “to see the data a bit better.” For example, in Task 3, P3 easily found the "Unknown" genre problem in the descriptive statistics, immediately updating the corresponding assumptions to handle it. Furthermore, P9 justified the usefulness of the aggregated information per columns by stating “it is something a lot of times you would end up asking about anyways,” and P1 appreciated that it “gives you a preview of everything together.”
@@ 6.3.7. Deferring Steering after Seeing Initial Results
Participants frequently preferred to first see results before interacting with and steering the AI tool. P6 highlighted that they “just want to see the result first and then trim it”. P18 wanted to “look at the result first and if the result was nonsense then go back”. This was particularly the case in the Phasewise and Stepwise systems, as P4 mentioned that the system made them “go through all of these steps and spend so much time before [they] could actually see the result.” P8 wished to see the generated code because they did not trust the system to generate correct code even if the execution plan was correct: “part of me wants to just see what it does, even if the [execution] plan looks reasonable, I know there’s gonna probably be errors.” Another reason why participants wanted to defer steering, particularly in the Phasewise system, was that the input/output assumptions or the execution plan did not have enough details about how the AI is going to “handle” or “calculate” something (P10, P14). Similarly, P12 was not sure about something in the plan, so they said “I’m going to generate first and see what happens.” Lastly, P4 indicated that having a small working memory was why they preferred the baseline system, where every interaction with the AI resulted in a complete output to verify.
@@ 6.4. Summary of Results
Our study revealed that while there was no difference in task success, completion time, or number of required verification hints, participants felt significantly more in control of the data analysis process when using the Phasewise and Stepwise systems compared to the Conversational baseline. Overall, the results show that both systems were preferred over the Conversational baseline. However, because the Phasewise system led to significantly higher information overload, the Stepwise system emerged as the most balanced and effective of the three.
@@ 7. Discussion and Implications for AI-Assisted Data Analysis Tools
  • DQ2 Steering Support: How does the user verify the current state of the AI’s output and determine which direction they should steer it? How should the tool facilitate users in making informed decisions at each steering point?
@@ 7.1. DQ1: Steering Points
Our results comparing user experiences with these systems reveals significant trade-offs. More frequent steering points increase users’ confidence in the results and their sense of control over the AI’s process. However, it demands more cognitive effort and delays the final outcome due to the frequent pauses created by the intervention points. Additionally, it may lead to premature decisions as users commit to directions without understanding their future impact. According to the Cognitive Dimensions of Notations framework (Green, 1989), this is a clear case of the system imposing “premature commitment”. Indeed, our findings indicate varied user preferences between the Conversational and Stepwise systems due to these factors.
@@ 7.2. DQ2: Steering Support
Participants noted their desire for the AI to act as an agent, aiding in the assumption-building process at steering points. Future tools could automatically retrieve assumptions by pinpointing specific, relevant evidence to support informed decision-making without overwhelming users. Additionally, these tools could help in accessing domain-specific knowledge pertinent to the data analysis task. Lastly, tools should be transparent regarding the AI’s limitations in sourcing all necessary information for optimal decision-making at each decision point.
@@ 7.3. DQ3: Steering Modality
After the user has decided which direction to steer the AI, their next action is to specify their intent to the AI. The design question here is determining the right interface and modality for the user to steer the AI. In response, we introduced two distinct interfaces: a free-form text editor used in the Conversational system for maximal flexibility, and structured editors in the Stepwise and Phasewise systems. The structured editors contain the assumptions and actions generated by the LLM’s chain-of-thought, allowing users to edit them for steering.
@@ 8.2. System Limitations
Another limitation of the system is that edits to downstream code or assumptions are not propagated to upstream assumptions, and if a user makes an earlier edit it will overwrite any subsequent changes. The intention behind this prima facie design decision was to present a simple model of “cause and effect” that represented how completions were generated, namely, the context of any point in the system is only that which appears before. Some participants identified this limitation and it influenced their steering preferences, choosing not to edit an assumption because it would clear changes that had been made to code.
@@ 9. Conclusion
In this work, we explore the design space of AI-assisted data analysis tools by presenting two novel interfaces that aim to improve steering and verification. Starting from the observation that task decomposition is an emerging characteristic of recent LLM-based systems, we developed two systems that explore different modes of interactive task decomposition, each based on unique trade-offs. The first, Stepwise, decomposes the problem step by step; the second, Phasewise, decomposes the problem into logical phases. Our evaluation demonstrates that users experienced a greater sense of control and confidence with our systems in comparison to a chat-based baseline. Still, task decomposition is not without preference or cost. Some users prefer to work through the task incrementally, whilst others prefer to see the plan upfront. Additionally, highly-structured decomposition can introduce cognitive burden. Consequently, we imagine that future AI interfaces will need to support adaptive decomposition that reacts to the user and task.